OpenStudy (cake):

What is the solution to the equation x/4 = 6/12? x= 1 x= 2 x= 4 x= 8

OpenStudy (anonymous):

x=6/12.4 :)

OpenStudy (anonymous):

@cake \[\frac{ x }{ 4 }=\frac{ 6 }{ 12 }\] Cross multiply: \[12x=24\] Divide both sides by 18 and you'll get x=2

OpenStudy (anonymous):

isn't \(\frac{6}{12}\) the same as \(\frac{1}{2}\) ?

OpenStudy (anonymous):

if \(\frac{x}{4}=\frac{1}{2}\) then \(x=4\times \frac{1}{2}=2\)

OpenStudy (cake):

2 or 6

OpenStudy (cake):

i guess 2

OpenStudy (texaschic101):

I agree with @SweetsAndEverything ...set it up as a proportion and then cross multiply,solving for x. Answer is 2.

OpenStudy (anonymous):

@cake Yes, it's 2. :)